# Federated Distributed Key Generation

### v2

@@ VI Liveness Simulations
To evaluate FDKG’s resilience under realistic deployment conditions, we conducted simulations exploring its performance under various parameter settings, comparing its liveness properties against expectations from traditional DKG under similar failure models. Our analysis focuses on three aspects: 1) parameter sensitivity, 2) scalability across network sizes, and 3) topology-dependent performance characteristics. The primary objective was to determine which configurations optimize liveness under varying conditions, thereby offering insights for practical deployment.
VI-A Simulation Methodology
We developed a simulation framework implementing both FDKG and standard DKG dynamics across two network topologies. Each simulation involves the following steps:
  1. 1.
    Network Generation: We considered two network topologies to represent different communication patterns:
    • Barabási-Albert (BA): A scale-free model designed to simulate scenarios where nodes exhibit preferential attachment, creating highly connected hubs.
    • Random Graph: A uniform model representing scenarios where nodes interact uniformly, with no highly connected hubs.
    Networks were initialized with a specified number of nodes ($n$). Guardian sets of size $k$ were then selected for participating nodes based on the specific topology rules (preferential attachment for BA, uniform for Random Graph).
  2. 2.
    Participant Selection: For each simulation run, subsets of nodes were selected to participate in the protocol phases:
    • Distribution Participants: A fraction $p$ of nodes were selected to participate, using either uniform random sampling (for the Random Graph model) or degree-weighted sampling (for the Barabási-Albert model). Let this set be $\mathbb{D}$.
    • Reconstruction Participants: A fraction $r$ of the initial Distribution Participants (from $\mathbb{D}$) were randomly selected to be available for the Reconstruction Phase, representing the retention rate. Let this set be $\mathbb{T}$.
  3. 3.
    Liveness Check: For each simulation run, the protocol checked whether the set of available reconstruction participants ($\mathbb{T}$) could successfully meet the conditions required for reconstructing the secret, respecting the FDKG rules. Success was defined as achieving Liveness of Reconstruction (Definition 1) despite node failures or absences modelled by the selection process above. Specifically, for each $P_{i}\in\mathbb{D}$, we checked if either $P_{i}\in\mathbb{T}$ or if at least $t$ guardians from $\mathcal{G}_{i}$ were present in $\mathbb{T}$. If this condition held for all $i\in\mathbb{D}$, the run was marked as successful.
@@ VI-A Simulation Methodology
The web simulator is available at: https://fdkg.stan.bar, source code is available at https://github.com/stanbar/fdkg. Figure 4 presents a snapshot of two different network topologies for the same setting.
Figure 4: Comparison of Random (left) and Barabási-Albert (right) guardian set selection policies for $n=100,p=0.3,r=0.9,k=5,t=3$. Gray nodes indicate the absence of the node, green nodes represent those that were present in both the distribution and reconstruction phases, blue nodes were present only in the first (distribution) phase, and orange nodes only in the second (reconstruction) phase.
VI-B Parameter Space and Simulation Workflow
We explored a comprehensive range of parameters to capture diverse deployment scenarios:
  • Number of Participants ($n$): $\{50,100,200,\dots,1000\}$.
  • Participation Rate ($p$): $\{0,1,0.2,\dots,0.9,1.0\}$.
  • Retention Rate ($r$): $\{0.1,0.2,\dots,1.0\}$.
  • Number of Guardians ($k$): $\{1,3,\dots,n-1\}$.
  • Threshold ($t$): Takes values from $1$ up to $k$.
For each unique combination of parameters, 100 independent simulations were executed, resulting in over 100 million total runs.
Figure 5 shows the scalability comparison between network models with fixed $p=1.0$, and $r=0.7$. The Y axis (success rate) is a mean value among all the $(k,t)$ configurations. Standard DKG corresponds to the FDKG case where $p=1.0$ and k=n1k=n-1 (all other nodes are guardians), thus its performance represents an upper bound on the mean success rate for a given $r$. In standard DKG each node connects to every other node, creating a complete graph, so there is no difference between BA and Random guardian selection policies. The BA policy (solid lines) generally outperforms the random policy (dashed lines) due to hub resilience, especially at smaller $n$.
Figure 5: Scalability comparison between average guardian set size and different policy selection with fixed $p=1.0,r=0.7$. The blue line represents the standard DKG scheme (k=n1k=n-1).
We then break down the mean values to see how each parameter influences the success rate. To focus on the impact of $p,r,k,t$, we subsequently use only the results from the BA policy selection.
Figure 6 shows success rates for each of the $(k,t)$ configurations for fixed $n=100$, $p=0.8$. Figure 6(a) uses $r=0.5$, while Figure 6(b) uses $r=0.9$, highlighting the critical significance of the retention parameter $r$.
Figure 6: Heatmaps of success rates by guardian configuration for fixed $n=100$, $p=0.8$, $r=0.5$ (top) and $r=0.9$ (bottom). Color intensity represents success probability from 0 (yellow) to 1 (dark blue).
Figure 7 reveals three critical relationships through our multivariate analysis. A higher retention rate ($r=0.9$) guarantees a high success rate even for a smaller number of guardians ($k=10$) at a partial participation rate ($p=0.75$). A lower retention rate ($r=0.5$) may require a higher participation rate ($\geq 0.6$) and a higher number of guardians ($k\geq 40$) to achieve a perfect success rate. Lower retention rates ($\leq 0.5$) combined with a low number of guardians ($k\leq 10$) may not guarantee high success rates even at perfect participation ($p=1.0$).
@@ VI-B Parameter Space and Simulation Workflow
Figure 7: Success rate vs participation rate (top) and retention rate (bottom) for varying $k$, and $p$, or $r$ combinations, with fixed $n=100$ and $t=k/4$. Line styles represent retention or participation rates, colors indicate guardian set sizes.
VI-C Results
The results reveal clear trends in the factors influencing the success rate (liveness probability). The participation rate ($p$) has a significant impact, with higher participation rates consistently improving success rates. This is because more nodes actively contributing in Round 1 increases the likelihood that enough participants or their guardians are available for reconstruction.
Retention rate ($r$) also plays a crucial role. Higher retention rates both reduce the reliance on guardians and increase the chances of meeting the threshold during reconstruction phase. For instance, when $r=1.0$, the protocol always achieves Liveness even with very low participation $p$ and guardians $k$, it is because the party who contributed to the first phase always reconstruct the share in the second phase. However, such scenarios were excluded from the analysis since they do not represent the challenges of FDKG systems.
The number of guardians ($k$) contributes to the robustness of the system, especially in scenarios with a larger total number of nodes. Larger guardian sets provide greater redundancy, mitigating the impact of absent nodes during reconstruction. This becomes increasingly important as the system scales, where the probability of absences rises proportionally with the total number of participants.
Threshold values ($t$) directly affect liveness. Lower thresholds (requiring fewer available guardians) are naturally more likely to succeed. However, higher thresholds can still succeed if combined with sufficiently large guardian sets ($k$), high participation ($p$), and high retention ($r$).
The results highlight also the impact of guardians selection policy on liveness, particularly the advantage of the Barabási-Albert model over the Random model. The superior performance of the Barabási-Albert model is primarily due to its weighted selection policy. In this policy, nodes with higher connectivity (hubs) are more likely to be selected as participants in both the Distribution and Reconstruction phases. This approach mirrors real-world social structures where highly connected individuals or entities play a more central role in communication and collaboration. Consequently, these hubs are not only more resilient to failures, but are also more likely to engage in critical democratic processes, such as voting or key reconstruction. However, the results show that with the higher network sizes the difference between those models diminishes.
To assess FDKG’s practical performance under realistic deployment conditions—such as node unavailability, varying participation rates, and network churn—we conducted simulations. These simulations complement the formal security guarantees provided in Section V, which ensure resilience against adversarial attacks under cryptographic assumptions. Importantly, the simulations do not aim to prove security; rather, they illustrate how protocol parameters like participation rate ($p$), retention rate ($r$), guardian set size ($k$), and threshold ($t$) impact the likelihood of successful key reconstruction in scenarios with non-adversarial failures (e.g., honest node crashes or dropouts). The results offer practical insights for parameter selection in real-world deployments, where both adversarial and non-adversarial challenges must be considered.
@@ VII-B Round 2: Vote Casting
    Vote Encryption (ElGamal): Generate a random blinding factor $r_{i}\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\mathbb{Z}_{q}$. Compute the encrypted ballot $B_{i}=(C1_{i},C2_{i})$ where:
@@
    Proof Generation: Generate a NIZK proof $\pi_{Ballot_{i}}$ demonstrating that $B_{i}$ is a valid ElGamal encryption of a vote $v_{i}$ from the allowed set $\{G^{2^{0}},G^{2^{m}},\dots,G^{2^{(c-1)m}}\}$ under public key E, without revealing $v_{i}$ or $r_{i}$. (See Appendix A-D).
@@ VII-C1 Online Tally
    • Generate a NIZK proof $\pi_{PD_{i}}$ proving that $\mathrm{PD}_{i}$ was computed correctly using the secret $d_{i}$ corresponding to the public $E_{i}$. (See Appendix A-E).
@@
    • Decrypt the share: $s_{j,i}=\texttt{Dec}_{\textrm{sk}_{i}}(C_{j,i})$.
@@
    • Generate a NIZK proof $\pi_{PDS_{j,i}}$ proving correct decryption of $C_{j,i}$ to $s_{j,i}$ and correct exponentiation of $C1$ by $s_{j,i}$. (See Appendix A-F).
@@ VIII Performance Evaluation
VIII-A Computational and Communication Costs per Message Type
Table II summarizes the core performance metrics for the different messages broadcast during the online phases of the protocol. It includes the time to generate the required NIZK proof and the total message size, which comprises the cryptographic payload plus the 256-byte Groth16 proof.
Table II: Performance Summary: NIZK Proving Time and Total Message Size per Broadcast Message Type.
Measurement FDKG Distribution Ballot Broadcast Partial Dec. Broadcast Part. Dec. Share Broadcast
@@ VIII-A Computational and Communication Costs per Message Type
The online computational cost is dominated by the generation of the $\pi_{FDKG_{i}}$ proof, especially as $t$ and $k$ increase. The communication cost for the FDKG distribution phase is primarily driven by $k$. Messages related to voting and tallying are significantly smaller individually, but the total communication cost of tallying depends heavily on how many secrets and shares need to be revealed publicly.
VIII-B Offline Tally: Vote Extraction Cost
@@
VIII-C Example Scenario Calculation
@@
  • $|\mathbb{D}|=50$ valid participants in Round 1 (FDKG Distribution).
@@
Using the ”Total Message Size (Bytes)” figures from Table II:
@@ X Limitations and Future Work
Guardian Set Selection: Our liveness simulations (Section VI) explored heuristics like Barabási-Albert for guardian selection, which showed promise. However, a deeper investigation into context-aware guardian selection strategies could further bolster FDKG’s practical reliability. This might involve incorporating social trust metrics, node reputation systems, or application-specific criteria to optimize for liveness and security against targeted attacks.
@@
Overall, these points highlight opportunities for optimizing cryptographic components, enhancing theoretical models (e.g., for adaptive adversaries in the FDKG context), and broadening the practical deployment of FDKG.
@@ XI Discussion and Conclusions
FDKG generalizes traditional DKG by integrating principles of federated trust, reminiscent of Federated Byzantine Agreement (FBA) systems like the Stellar Consensus Protocol. This allows for dynamic trust delegation at the participant level. By achieving its core functionality in two communication rounds (distribution and reconstruction), FDKG is suitable for systems involving human interaction, such as the i-voting application presented. The reliance on NIZK proofs (e.g., zk-SNARKs) for verifiability ensures protocol integrity but introduces computational overhead and, in the case of Groth16, a trusted setup requirement, which are important considerations for deployment. Future work could explore NIZKs with transparent setups (e.g., STARKs or Bulletproofs, though the latter might impact succinctness) to alleviate this.
@@
Performance evaluations (Section VIII) quantified the computational and communication costs. For $n=100$ participants with $k=40$ guardians, the FDKG distribution phase incurred approximately 336 kB of broadcast data per participant. NIZK proof generation, while enabling verifiability, is the main computational bottleneck, with $\pi_{FDKG_{i}}$ proving times ranging from $3.36$s (for $k=10$) to $14.79$s (for $k=100$) in our experiments (Table II). These figures highlight the trade-off between the flexibility offered by larger guardian sets and the associated costs, underscoring the need for careful parameter tuning based on specific application constraints: smaller $(k,t)$ values may be preferable for latency-sensitive systems, while larger values enhance resilience in more adversarial settings. A critical aspect of FDKG is balancing guardian set size ($k$) against security; while smaller sets reduce overhead, they concentrate trust, making individual guardians more critical for liveness if the original participant is unavailable or malicious.
We envision FDKG as a versatile cryptographic primitive for multi-party systems that require robust and verifiable key distribution but must operate under conditions of incomplete or dynamic participation. Future research directions include optimizing NIZK proofs for FDKG-specific relations, refining vote decoding mechanisms for FDKG-based voting schemes, and developing more sophisticated, context-aware guardian selection strategies to further enhance FDKG’s practical usability and security. By embracing a decentralized and adaptive ethos, FDKG offers a pathway to building scalable and secure distributed systems for diverse collaborative environments.
@@
To ensure the reproducibility of our findings, all source code, datasets, and scripts used to generate results are publicly available. The repository is hosted at https://github.com/stanbar/fdkg.
Acknowledgment

### v3

@@ VI Liveness Simulations
This section empirically evaluates liveness under non-adversarial (random) churn. The goal is to estimate, for a configuration $(n,p,r,k,t)$ and a guardian-selection policy, the probability that the Round 2 reconstruction predicate induced by (1) is satisfied. These experiments complement the conditions in Theorems 4 and 3: they do not constitute a cryptographic proof, but quantify the operational regimes where reconstruction is likely to succeed in the presence of availability failures.
Evaluation predicate
Let $\mathbb{D}$ be the set of valid Round 1 participants and $\mathbb{T}\subseteq\mathbb{P}$ be the Round 2 actors that are present. A trial is a success iff
$\forall i\in\mathbb{D}:\quad\bigl(i\in\mathbb{T}\bigr)\ \ \lor\ \ \bigl(\bigl|\mathbb{T}\cap\mathcal{G}_{i}\bigr|\geq t\bigr),$ (2)
i.e., iff $\mathbb{T}\in R$ with $R$ as in (1); see Theorem 4.
VI-A Experimental Design and Rationale
Topology policies (guardian selection)
We study two canonical policies that bracket real deployments:
  • Random (Erdős–Rényi, ER): guardians are chosen uniformly; this models hub-free environments (uniform lists, randomized assignment).
  • Barabási–Albert (BA): guardians are chosen with preferential attachment; this models environments with socially/operationally prominent “hubs” (recognizable, responsive, institutionally visible parties).
Both policies enforce $|\mathcal{G}_{i}|=k$ with local threshold $t\leq k$. ER provides a neutral baseline without hubs; BA captures heavy-tailed guardian popularity, which concentrates redundancy around hubs and (under random churn) tends to increase the number of disjoint reconstruction paths.
Parameters and sampling
We explore $n\in\{50,100,200,\dots,1000\}$, $p\in\{0.1,0.2,\dots,1.0\}$, $r\in\{0.1,0.2,\dots,1.0\}$, $k\in\{1,3,\dots,n-1\}$, and $t\in\{1,\dots,k\}$. Round 1 participation draws $\mathbb{D}$ with $|\mathbb{D}|\approx p\,n$. Round 2 availability draws $\mathbb{T}$ with expected size $\approx r\,|\mathbb{P}|$. Unless stated otherwise, we run 100 independent, identically distributed (i.i.d.) trials per configuration and report the success rate (fraction of trials satisfying (2)).
DKG vs FDKG
Classical $(t,n)$-DKG is the special case of FDKG with full participation and maximal guardians:
$\text{DKG}(t,n)\equiv\text{FDKG}\bigl(p{=}1,\ t,\ k{=}n{-}1\bigr).$
Therefore, any comparison at identical parameters is tautological. In figures we include a reference line labelled “DKG” defined as the FDKG curve at (p=1,k=n1)(p{=}1,\ k{=}n{-}1); all other curves are FDKG with $(p\leq 1,\ k\leq n{-}1)$. This makes explicit that differences arise from configuration, not from different protocols.
Implementation and artifacts
@@ VI-A Experimental Design and Rationale
We use the public simulator at https://fdkg.stan.bar (source: https://github.com/stanbar/fdkg). Figure 3 illustrates ER vs. BA for the same $(n,p,r,k,t)$.
Figure 3: ER (left) vs. BA (right) guardian-selection examples for $n=100,p=0.3,r=0.9,k=5,t=3$. Gray: absent; green: present in both rounds; blue: Round 1 only; orange: Round 2 only.
@@
Figure 4: Success vs. participation $p$ (top) and retention $r$ (bottom) at $n=100$ with $t=k/4$. Line styles encode the other axis; colors encode $k$.
Figure 5: Success over $(k,t)$ for $n=100$, $p=0.8$, with $r=0.5$ (top) and $r=0.9$ (bottom). Darker indicates higher success. Higher $r$ expands the feasible region (larger $t$ at fixed $k$).
Figure 6: Scalability vs. guardian policy for p=1.0,r=0.7p=1.0,\ r=0.7. The DKG line is the FDKG configuration (p=1,k=n1)(p{=}1,\ k{=}n{-}1) (classical $(t,n)$-DKG). BA (solid) generally exceeds ER (dashed) under random churn; the gap narrows with larger $n$.
VI-B Results and Interpretation
Retention dominates
Retention $r$ is the primary driver of liveness. With $r\approx 0.8$, high success is attainable even for modest $k$ at partial participation. For instance, for $n=100$, $p=0.7$, $k=10$ achieves high success. Increasng $k=40$ yields near-certainty even at $p=0.5$ (Figure 4).
With $r\approx 0.5$, approaching certainty requires both higher $p$ and larger $k$ (e.g., for $n=100$, $p\gtrsim 0.7$ and $k\gtrsim 40$) (Figure 4).
Threshold vs. redundancy
For fixed $(n,p,r,k)$, success decreases as $t/k$ increases. Empirically, for $n=100$, $p=0.8$, $r=0.9$, liveness remains robust up to roughly $t\lesssim 0.7k$; with $r=0.5$ robustness holds up to roughly $t\lesssim 0.4k$ (Figure 5).
Topology under random churn
BA generally outperforms ER for small/medium $n$ because preferential attachment concentrates redundancy around hubs, creating multiple disjoint reconstruction paths. As $n$ grows, the gap narrows (Figure 6). This advantage pertains to random availability failures and does not imply robustness to targeted hub failures.
Scaling envelope
Across $n$, the DKG line (FDKG at p=1p{=}1, k=n1k{=}n{-}1) acts as an upper envelope at a given $r$. FDKG approaches this envelope as $k$ increases; for smaller $k$, it trades communication cost for liveness probability in a topology- and $r$-dependent manner.
VI-C Adaptive Adversaries
The security analysis (Section V) characterizes reconstruction privacy via $R$ (Theorem 3) and liveness via the existence of an honest $S\in R$ (Theorem 4). Under adaptive corruption at Round 2, two objectives diverge:
Targeting liveness
To falsify (2) for $i$, one must simultaneously prevent direct revelation (corrupt $i$) and reduce the remaining honest guardians below $t$ (thus corrupt $\geq k{-}t{+}1$ guardians).
Targeting privacy
Privacy fails iff the corrupted set contains a reconstruction-capable set for all participants’ guardian requirements (Theorem 3), i.e., there exists $S\subseteq\mathcal{C}$ with $S\in R$. The attacker needs to control enough guardians to reconstruct every participant’s partial secret (unless that participant is directly corrupted). How many corruptions are needed depends mainly on how guardian sets overlap:
  • Full overlap ($G_{i}=G$ for all $i$): corrupt any $t$ in $G$ to reconstruct all partial secrets (recovers the classical DKG threshold).
  • Disjoint guardians ($G_{i}\cap G_{j}=\varnothing$): the adversary must, for each $i$, either corrupt $i$ or $t$ members of $G_{i}$, so the optimum scales with the number of participants unless many $i$ are directly corrupted.
Thus, diverse guardian choices improve availability under random churn (Section VI-B), but privacy under targeted attacks depends on guardian-set overlaps. Operationally, implementations should provide UIs that encourage honest parties to choose diverse guardian sets to avoid over-concentration.
@@ VII-B Round 2: Vote Casting
    Vote Encryption (ElGamal): Generate a random blinding factor $r_{i}\>{\leftarrow\!\!\mbox{\tiny${\mathdollar}$\normalsize}}\,\mathbb{Z}_{q}$. Compute the encrypted ballot $B_{i}=(C1_{i},C2_{i})$ where:
@@
    Proof Generation: Generate a NIZK proof $\pi_{Ballot_{i}}$ demonstrating that $B_{i}$ is a valid ElGamal encryption of a vote $v_{i}$ from the allowed set $\{G^{2^{0}},G^{2^{m}},\dots,G^{2^{(c-1)m}}\}$ under public key E, without revealing $v_{i}$ or $r_{i}$. (See Appendix B-D).
@@ VII-C1 Online Tally
    • Generate a NIZK proof $\pi_{PD_{i}}$ proving that $\mathrm{PD}_{i}$ was computed correctly using the secret $d_{i}$ corresponding to the public $E_{i}$. (See Appendix B-E).
@@
    • Decrypt the share: $s_{j,i}=\textrm{Dec}_{\textrm{sk}_{i}}(C_{j,i})$.
@@
    • Generate a NIZK proof $\pi_{PDS_{j,i}}$ proving correct decryption of $C_{j,i}$ to $s_{j,i}$ and correct exponentiation of $C1$ by $s_{j,i}$. (See Appendix B-F).
@@ VIII Performance Evaluation
Computational and Communication Costs per Message Type
Table III summarizes the core performance metrics for the different messages broadcast during the online phases of the protocol. It includes the time to generate the required NIZK proof and the total message size, which comprises the cryptographic payload plus the 256-byte Groth16 proof.
Table III: Performance Summary: NIZK Proving Time and Total Message Size per Broadcast Message Type.
Measurement FDKG Generation Ballot Broadcast Partial Dec. Broadcast Part. Dec. Share Broadcast
@@
The online computational cost is dominated by the generation of the $\pi_{FDKG_{i}}$ proof, especially as $t$ and $k$ increase. The communication cost for the FDKG Generation phase is primarily driven by $k$. Messages related to voting and tallying are significantly smaller individually, but the total communication cost of tallying depends heavily on how many secrets and shares need to be revealed publicly.
Offline Tally: Vote Extraction Cost
@@
Example Scenario Calculation
@@
  • $|\mathbb{D}|=50$ valid participants in Round 1 (FDKG Generation).
@@
Using the ”Total Message Size (Bytes)” figures from Table III:
@@
Large-scale scenario ($n=5000$)
Assume $|\mathbb{D}|=2500$, $k=40$, $|\mathbb{V}|=2500$, $|\mathbb{T}\cap\mathbb{D}|=2000$, and $40$ shares revealed per tallier (80,00080{,}000 shares total). Using Table III sizes:
[…]
Communication is dominated by share revelations; for fixed $k$ and average shares per tallier, growth is near-linear in $|\mathbb{D}|$.
Scalability and load balance
FDKG uses one broadcast round in Generation and one in Reconstruction. End-to-end latency is dominated by (i) local proving time for $\pi_{FDKG_{i}}$ and (ii) network propagation per round; Reconstruction latency additionally depends on the number of shares each online guardian reveals and proves. Load is uneven by design: some guardians receive more delegations and therefore handle more decryption-share work. Let $c_{g}$ denote the number of parties that selected guardian $g$ (guardian “popularity”). A simple way to summarize skew is to report $\max_{g}c_{g}$, the median of $\{c_{g}\}$, and the ratio $\max_{g}c_{g}/\mathrm{median}(\{c_{g}\})$. In our experiments, this ratio is larger under BA than ER (reflecting hubs). Operators can bound peak load by capping per-guardian assignments, randomizing a fraction of guardian picks, or recommending diversified guardian sets.
The Generation broadcast cost scales as $\mathcal{O}(n\cdot k)$ elements (Section IV); with fixed $k$ the total cost is linear in $n$. Reconstruction can reach $\mathcal{O}(n^{2})$ in worst-case guardian topologies (some parties chosen by all other participants), but in practice remains near $\mathcal{O}(n\cdot k)$ when overlap is moderate. Hierarchical deployments (partitioning parties into domains, then aggregating domain-level partial keys) can reduce the maximum number of counterparties any single party must send to or serve during Generation/Reconstruction, while preserving heterogeneous trust.
@@ X Limitations and Future Work
Usability and guardian selection: Non-experts may find manual guardian selection burdensome. In practice, UIs can (i) offer curated recommendations that blend centrality (availability) with diversity (to avoid over-concentration), (ii) enforce soft caps on per-guardian assignments, (iii) set $(k,t)$ defaults from observed retention (e.g., $t\approx 0.3$$0.5$ of $k$ for moderate $r$), and (iv) warn about privacy risks under highly overlapping guardian choices. These controls preserve FDKG’s heterogeneous trust while reducing user burden.
@@
Resharing and forward secrecy: FDKG, as presented, is a bootstrapping primitive. Enabling long-lived deployments with proactive defense against key accumulation requires a dynamic resharing protocol that (i) changes guardian sets $G_{i}$ and/or shares without re-running Generation, and (ii) provides forward secrecy against state compromise. Integrating proactive secret sharing (DPSS) with heterogeneous guardian sets while preserving public verifiability and succinct proofs is an open direction.
Adaptive and targeted adversaries: So far the analysis assumes static corruption and random churn. Future work might study adaptive/rushing behavior and targeted removals against the guardian topology $\{G_{i}\}$. The goal is to quantify how privacy and liveness degrade when the attacker selects corruptions after seeing Round 1 transcripts, and to assess practical mitigations: mild commit–reveal or public randomness to limit key bias, plus topology-aware defenses. Using the same reconstruction predicate $R$, one could report topology-sensitive lower bounds on the effort needed to break privacy or liveness and identify when targeted strategies matter in practice.
Topology and load control: Hotspots and concentration risks can be reduced by combining local diversity constraints (soft/hard caps per guardian, randomly sampled guardians, overlap-distance limits, reputation that discounts hubs) with a hierarchical FDKG where domains run local FDKG and a second tier aggregates domain keys. Future work might measure how these knobs reshape $\{G_{i}\}$: lowering $\max_{g}c_{g}/\mathrm{median}(c_{g})$, bounding worst-case reconstruction traffic, and increasing the minimal reconstruction-capable set for targeted breaks—while preserving heterogeneous trust and near $O(nk)$ per tier.
@@ XI Discussion and Conclusions
FDKG generalizes traditional DKG by integrating principles of federated trust, reminiscent of Federated Byzantine Agreement (FBA) systems like the Stellar Consensus Protocol. This allows for dynamic trust delegation at the participant level. By achieving its core functionality in two communication rounds (generation and reconstruction), FDKG is suitable for systems involving human interaction, such as the i-voting application presented. The reliance on NIZK proofs (e.g., zk-SNARKs) for verifiability ensures protocol integrity but introduces computational overhead and, in the case of Groth16, a trusted setup requirement, which are important considerations for deployment. Future work could explore NIZKs with transparent setups (e.g., STARKs or Bulletproofs, though the latter might impact succinctness) to alleviate this.
@@
Performance evaluations (Section VIII) quantified the computational and communication costs. For $n=100$ participants with $k=40$ guardians, the FDKG Generation phase incurred approximately 336 kB of broadcast data per participant. NIZK proof generation, while enabling verifiability, is the main computational bottleneck, with $\pi_{FDKG_{i}}$ proving times ranging from $3.36$s (for $k=10$) to $14.79$s (for $k=100$) in our experiments (Table III). These figures highlight the trade-off between the flexibility offered by larger guardian sets and the associated costs, underscoring the need for careful parameter tuning based on specific application constraints: smaller $(k,t)$ values may be preferable for latency-sensitive systems, while larger values enhance resilience in more adversarial settings. A critical aspect of FDKG is balancing guardian set size ($k$) against security; while smaller sets reduce overhead, they concentrate trust, making individual guardians more critical for liveness if the original participant is unavailable or malicious.
We envision FDKG as a versatile cryptographic primitive for multi-party systems that require robust and verifiable key generation but must operate under conditions of incomplete or dynamic participation. Future research directions include optimizing NIZK proofs for FDKG-specific relations, refining vote decoding mechanisms for FDKG-based voting schemes, and developing more sophisticated, context-aware guardian selection strategies to further enhance FDKG’s practical usability and security. By embracing a decentralized and adaptive ethos, FDKG offers a pathway to building scalable and secure distributed systems for diverse collaborative environments.
@@
To ensure the reproducibility of our research, all source code, datasets, and scripts used to generate results are publicly available. The repository is hosted at https://github.com/stanbar/fdkg.
Acknowledgments